# On denominators of consecutive $\operatorname{SL}(2,\mathbb{N})$-saturated Farey fractions

### v1

@@ 3. The inclusion $\mathcal{V}\subseteq\overline{\mathscr{D}}$
This shows that we need to find at least one integer triple $(q_{1},q_{2},b)$ such that [^1] Here, $b$ plays the role of $\bar{q}_{2}=q_{1}-a_{1}$.
$q_{1}\in QI_{1},\quad q_{2}\in QI_{2},\quad q_{2}b\equiv-1\left(\mathrm{mod\;}q_{1}\right),\quad 3q_{1}-q_{2}-Q<b<q_{1}.$
The inequalities $\max\{Q-q_{1},2q_{1}-Q\}<q_{2}<q_{1}$ are plain consequences of the inclusion $\mathcal{R}=I_{1}\times I_{2}\subseteq\mathring{\mathcal{V}}_{1}$. The set of such integer triples is given by
$\begin{split}\sum\limits_{q_{1}\in QI_{1}}&\bigg{(}\frac{\varphi(q_{1})}{q_{1}^{2}}\operatorname{Area}\big{\{}(q_{2},b):q_{2}\in QI_{2},3q_{1}-q_{2}-Q<b<q_{1}\big{\}}+O_{\varepsilon}(q^{1/2+\varepsilon})\bigg{)}\\ &=\frac{Q^{2}}{\zeta(2)}\int_{QI_{1}}\frac{dx}{x}\operatorname{Area}\big{\{}(y,z):y\in I_{2},3x-y-1<z<x\big{\}}+O_{\varepsilon}(Q^{3/2+\varepsilon}).\end{split}$
@@
$H_{1}(x,y)=\frac{1+y-2x}{x}>0\quad\text{ on $\mathring{\mathcal{V}}_{1}$},$ (8)
allowing us to conclude the proof.
We now look at the case where $x_{0}+y_{0}<1$. Although this case is similar to the previous one, it is a bit more involved since there will be one or possibly more fractions in $\mathcal{F}_{Q}$ between $a_{1}/q_{1}$ and $a_{2}/q_{2}$. In order to deal with these intermediate fractions, we use the following result:
@@ 2. $(2)$
Assume first that $(x_{0},y_{0})\in\mathring{\mathcal{V}}_{2}$, so $(x_{0},y_{0})\in\mathring{\mathcal{V}}$, $x_{0}+y_{0}<1$, and $x_{0}<y_{0}$. We prove that for large $Q$ there exist integer triples $(q_{1},q_{2},a)$ such that $q_{2}>q_{1}$, $q_{1}a\equiv 1\left(\mathrm{mod\;}q_{2}\right)$, and [^2] Here, $a$ plays the role of $\bar{\bar{q}}_{1}$.
$\begin{split}&(q_{1},q_{2})\in Q\mathcal{R},\quad 0<a<q_{2},\\ &\bigg{(}2+\bigg{\lfloor}\frac{q_{2}}{q_{1}}\bigg{\rfloor}\bigg{)}q_{1}-q_{2}+\frac{q_{1}a-1}{q_{2}}\leq Q,\quad(\Longleftrightarrow h(a_{1}/q_{1})\leq Q)\\ &q_{1}+q_{2}+a\leq Q,\quad(\Longleftrightarrow h(a_{2}/q_{2})\leq Q)\\ &2q_{1}+q_{2}+a+\frac{q_{1}a-1}{q_{2}}>Q.\quad(\Longleftrightarrow h(a_{*}/q_{*})>Q)\end{split}$ (10)
@@
Employing standard techniques, we find that the number of triples $(q_{1},q_{2},a)$ as in (10) is given by
$\frac{Q^{2}}{\zeta(2)}\iint_{\mathcal{R}}H_{2}(x,y)\,dxdy+O_{\varepsilon}(Q^{3/2+\varepsilon}),$
@@
$H_{2}(x,y)=\min\bigg{\{}\frac{1+y-(2+\lfloor y/x\rfloor)x}{x},\frac{1-x-y}{y},1\bigg{\}}-\max\bigg{\{}\frac{1-2x-y}{x+y},0\bigg{\}},$ (11)
with $H_{2}(x,y)>0$ on $\mathring{\mathcal{V}}_{2}$.
Assume next that $(x_{0},y_{0})\in\mathring{\mathcal{V}}_{3}$, so $(x_{0},y_{0})\in\mathcal{V}$, $x_{0}+y_{0}<1$ and $y_{0}<x_{0}$. We prove that for large $Q$ there exist integer triples $(q_{1},q_{2},b)$ such that $q_{2}<q_{1}$, $q_{2}b\equiv-1\left(\mathrm{mod\;}q_{1}\right)$, and [^3] Here, $b$ plays the role of $\bar{q}_{2}=q_{1}-a_{1}$.
$\begin{split}&(q_{1},q_{2})\in Q\mathcal{R},\quad 0<b<q_{1},\\ &3q_{1}-q_{2}-b\leq Q,\quad(\Longleftrightarrow h(a_{1}/q_{1})\leq Q)\\ &q_{1}+\bigg{(}2-\bigg{\lfloor}\frac{q_{1}}{q_{2}}\bigg{\rfloor}\bigg{)}q_{2}-\frac{q_{2}b-1}{q_{1}}\leq Q,\quad(\Longleftrightarrow h(a_{2}/q_{2})\leq Q)\\ &3q_{1}+2q_{2}-b-\frac{q_{2}b-1}{q_{1}}>Q.\quad(\Longleftrightarrow h(a_{*}/q_{*})>Q)\end{split}$ (12)
The inequalities in (12) are equivalent to
By a plain check, we see that
$F_{Q}(q_{1},q_{2})\leq\max\bigg{\{}3q_{1}-q_{2}-Q,\frac{1+q_{1}(3q_{2}-Q)}{q_{2}},0\bigg{\}}<G_{Q}(q_{1},q_{2}).$
Employing standard techniques, we find that the number of triples $(q_{1},q_{2},a)$ as in (12) is given by
$\frac{Q^{2}}{\zeta(2)}\iint_{\mathcal{R}}H_{3}(x,y)\,dxdy+O_{\varepsilon}(Q^{3/2+\varepsilon}),$
where
$H_{3}(x,y)=\min\bigg{\{}\frac{3x+2y-1}{x+y},1\bigg{\}}-\max\bigg{\{}\frac{3x-y-1}{x},\frac{x+(2-\lfloor x/y\rfloor)y-1}{y},0\bigg{\}},$ (13)
with $H_{3}(x,y)>0$ on $\mathring{\mathcal{V}}_{3}$.
@@ 4. A partition of the region $\mathcal{V}$
$\bigg{(}\frac{q_{1}}{Q},\frac{q_{2}}{Q}\bigg{)}\in\mathcal{W}_{1}:=\mathcal{V}_{1}\subseteq\bigg{[}0,\frac{1}{2}\bigg{]}\times\bigg{[}\frac{1}{3},1\bigg{]}.$
@@
$\bigg{(}\frac{q_{1}}{Q},\frac{q_{4}}{Q}\bigg{)}\in\mathcal{W}_{3,1}:=\bigg{\{}(x,y)\in[0,1]^{2}:\max\bigg{\{}1-2x,\frac{1-x}{3},2x-1\bigg{\}}\leq y\leq\frac{1-x}{2}\bigg{\}},$
with $\mathcal{W}_{3,1}\subseteq[1/3,3/5]\times[1/7,1/3]$.
@@
$\bigg{(}\frac{q_{1}}{Q},\frac{q_{4}}{Q}\bigg{)}\in\mathcal{W}_{3,2}:=\bigg{\{}(x,y)\in[0,1]^{2}:\max\bigg{\{}1-3x,\frac{1-x}{2}\bigg{\}}\leq y\leq 1-2x\bigg{\}},$
with $\mathcal{W}_{3,2}\subseteq[0,1/3]\times[1/3,1]$.
@@
$\bigg{(}\frac{q_{1}}{Q},\frac{q_{5}}{Q}\bigg{)}\in\mathcal{W}_{4,1}:=\bigg{\{}(x,y)\in[0,1]^{2}:\max\bigg{\{}1-2x,\frac{1-x}{4},2x-1\bigg{\}}\leq y\leq\frac{1-x}{3}\bigg{\}},$
with $\mathcal{W}_{4,1}\subseteq[2/5,4/7]\times[1/9,1/5]$.
@@
$\bigg{(}\frac{q_{1}}{Q},\frac{q_{5}}{Q}\bigg{)}\in\mathcal{W}_{4,2}:=\bigg{\{}(x,y)\in[0,1]^{2}:\max\bigg{\{}1-3x,\frac{1-x}{3}\bigg{\}}\leq y\leq\min\bigg{\{}\frac{1-x}{2},1-2x\bigg{\}}\bigg{\}},$
with $\mathcal{W}_{4,2}\subseteq[1/5,2/5]^{2}$.
@@
$\bigg{(}\frac{q_{1}}{Q},\frac{q_{5}}{Q}\bigg{)}\in\mathcal{W}_{4,3}:=\bigg{\{}(x,y)\in[0,1]^{2}:\max\bigg{\{}1-4x,\frac{1-x}{2}\bigg{\}}\leq y\leq 1-3x\bigg{\}},$
with $\mathcal{W}_{4,3}\subseteq[0,1/5]\times[2/5,1]$.
@@
$\bigg{(}\frac{q_{1}}{Q},\frac{q_{6}}{Q}\bigg{)}\in\mathcal{W}_{5,1}:=\bigg{\{}(x,y)\in[0,1]^{2}:\max\bigg{\{}1-2x,\frac{1-x}{5},2x-1\bigg{\}}\leq y\leq\frac{1-x}{4}\bigg{\}},$
with $\mathcal{W}_{5,1}\subseteq[3/7,6/11]\times[1/11,1/7]$.
@@
$\bigg{(}\frac{q_{1}}{Q},\frac{q_{6}}{Q}\bigg{)}\in\mathcal{W}_{5,2}:=\bigg{\{}(x,y)\in[0,1]^{2}:\max\bigg{\{}1-3x,\frac{1-x}{4}\bigg{\}}\leq y\leq\min\bigg{\{}\frac{1-x}{3},1-2x\bigg{\}}\bigg{\}},$
with $\mathcal{W}_{5,2}\subseteq[1/4,3/7]\times[1/7,1/4]$.
@@
$\bigg{(}\frac{q_{1}}{Q},\frac{q_{6}}{Q}\bigg{)}\in\mathcal{W}_{5,3}:=\bigg{\{}(x,y)\in[0,1]^{2}:\max\bigg{\{}1-4x,\frac{1-x}{3}\bigg{\}}\leq y\leq\min\bigg{\{}\frac{1-x}{2},1-3x\bigg{\}}\bigg{\}},$
with $\mathcal{W}_{5,3}\subseteq[1/7,1/4]\times[1/4,3/7]$.
@@
$\bigg{(}\frac{q_{1}}{Q},\frac{q_{6}}{Q}\bigg{)}\in\mathcal{W}_{5,4}:=\bigg{\{}(x,y)\in[0,1]^{2}:\max\bigg{\{}1-5x,\frac{1-x}{2}\bigg{\}}\leq y\leq 1-4x\bigg{\}},$
with $\mathcal{W}_{5,4}\subseteq[0,1/7]\times[3/7,1]$.
Next, we show the existence of examples where the boundary line $3x+2y=1$ of $\mathcal{V}$ occurs as a side of a region $\mathcal{W}_{r,i}$. The tuple $(q_{1},\ldots,q_{10})=(27,95,68,41,96,55,69,83,97,14)$ represents consecutive elements in $\mathcal{F}_{100}$ with $q_{1}=27$ and $q_{10}=84$ consecutive in $\mathscr{S}_{100}$ and mediant $q_{1}+q_{10}=41=q_{4}$. The equalities $q_{2}=3q_{1}+q_{10}$, $q_{3}=2q_{1}+q_{10}$, $q_{4}=q_{1}+q_{10}$, $q_{5}=2q_{1}+3q_{10}$, $q_{6}=q_{1}+2q_{10}$, $q_{7}=q_{1}+3q_{10}$, $q_{8}=q_{1}+4q_{10}$, $q_{9}=q_{1}+5q_{10}$ show that the region $Q\mathcal{W}_{9,4}$ is defined by the inequalities $\max\{3q_{1}+q_{10},2q_{1}+3q_{10},q_{1}+5q_{10}\}\leq Q<\min\{q_{1}+q_{2}=4q_{1}+q_{10},q_{3}+q_{4}=3q_{1}+2q_{10},q_{6}+q_{7}=2q_{1}+5q_{10},q_{9}+q_{10}=q_{1}+6q_{10}\}$, and thus

### v2

@@ 3. The inclusion $\mathcal{V}\subseteq\overline{\mathscr{D}}$ and the proof of Theorem 2
This shows that we need to count the integer triples $(q_{1},q_{2},b)$ such that [^1] Here, $b$ plays the role of $\bar{q}_{2}=q_{1}-a_{1}$. Note also that $3q_{1}-q_{2}-Q>2q_{1}-Q>0$.
$q_{1}\in QI_{1},\quad q_{2}\in QI_{2},\quad q_{2}b\equiv 1\left(\mathrm{mod\;}q_{1}\right),\quad 3q_{1}-q_{2}-Q\leq b<q_{1}.$
The inequalities $\max\{Q-q_{1},2q_{1}-Q\}<q_{2}<q_{1}$ follow from $\mathcal{R}=I_{1}\times I_{2}\subseteq\mathring{\mathcal{V}}_{1}$. Employing Lemma 2.3 in [6] and Lemma 3.1 in [5], the set of such integer triples is seen to be given by
$\begin{split}\sum\limits_{q_{1}\in QI_{1}}&\sum\limits_{\begin{subarray}{c}q_{2}\in QI_{2}\\ 3q_{1}-q_{2}-Q\leq b<q_{1}\\ q_{2}b\equiv 1\left(\mathrm{mod\;}q_{1}\right)\end{subarray}}1\\ &=\sum\limits_{q_{1}\in QI_{1}}\bigg(\frac{\varphi(q_{1})}{q_{1}^{2}}\operatorname{Area}\big\{(q_{2},b):q_{2}\in QI_{2},3q_{1}-q_{2}-Q\leq b\leq q_{1}\big\}+O_{\varepsilon}(q_{1}^{1/2+\varepsilon})\bigg)\\ &=\frac{Q^{2}}{\zeta(2)}\int_{I_{1}}\frac{dx}{x}\operatorname{Area}\big\{(y,z):y\in I_{2},3x-y-1\leq z\leq x\big\}+O_{\varepsilon}(Q^{3/2+\varepsilon}).\end{split}$
@@
$H_{1}(x,y)=\frac{1+y-2x}{x}>0\quad\text{ on $\mathring{\mathcal{V}}_{1}$}.$ (10)
In conjunction with (2), this allows us to conclude the proof.
We now look at the case where $x_{0}+y_{0}<1$. Although analogous to the previous one, this is slightly more involved since there will be one or possibly more fractions in $\mathcal{F}_{Q}$ between $a_{1}/q_{1}$ and $a_{2}/q_{2}$. In order to deal with these intermediate fractions, we use the following result:
@@ 2. $(2)$
Assume now that $\mathcal{R}\subseteq\mathring{\mathcal{V}}_{2}$, so $x_{0}+y_{0}<1$ , $x_{0}>y_{0}$, $3x_{0}+2y_{0}>1$ and $1+y_{0}-2x_{0}>0$. Here, the integer pairs $(q_{1},q_{2})\in\mathscr{D}_{Q}\cap Q\mathcal{R}$ must satisfy the inequalities $q_{1}+q_{2}<Q$, $q_{1}>q_{2}$ and $3q_{1}+2q_{2}>Q>2q_{1}-q_{2}$. Taking (11) into account, this is equivalent with counting integer triples $(q_{1},q_{2},b)$ such that [^2] Here, $b$ plays the role of $\bar{q}_{2}=q_{1}-a_{1}$.
$\begin{split}&q_{1}\in QI_{1},\quad q_{2}\in QI_{2},\quad q_{1}>q_{2},\quad 0<b<q_{1},\quad q_{2}b\equiv 1\left(\mathrm{mod\;}q_{1}\right),\\ &3q_{1}-q_{2}-b\leq Q,\quad(\Longleftrightarrow h(a_{1}/q_{1})\leq Q)\\ &q_{1}+\bigg(2-\bigg\lfloor\frac{q_{1}}{q_{2}}\bigg\rfloor\bigg)q_{2}-\frac{q_{2}b-1}{q_{1}}\leq Q,\quad(\Longleftrightarrow h(a_{2}/q_{2})\leq Q)\\ &3q_{1}+2q_{2}-b-\frac{q_{2}b-1}{q_{1}}>Q.\quad(\Longleftrightarrow h(a_{*}/q_{*})>Q)\end{split}$ (12)
The inequalities in (12) are equivalent to
By a plain check, we see that
$F_{Q}(q_{1},q_{2})\leq\max\bigg\{3q_{1}-q_{2}-Q,\frac{1+q_{1}(3q_{2}-Q)}{q_{2}},0\bigg\}\leq G_{Q}(q_{1},q_{2}).$
Employing the same summation technique as in $\mathcal{R}\subseteq\mathring{\mathcal{V}}_{1}$, the number of triples $(q_{1},q_{2},b)$ as in (12) is seen to be given by
$\sum\limits_{q_{1}\in QI_{1}}$ $\sum\limits_{\begin{subarray}{c}q_{2}\in QI_{2}\\ F_{Q}(q_{1},q_{2})\leq b\leq G_{Q}(q_{1},q_{2})\\ q_{2}b\equiv 1\left(\mathrm{mod\;}q_{1}\right)\end{subarray}}1$
$=\sum\limits_{q_{1}\in QI_{1}}\bigg(\frac{\varphi(q_{1})}{q_{1}^{2}}\operatorname{Area}\big\{(b,q_{2}):q_{2}\in QI_{2},F_{Q}(q_{1},q_{2})\leq b\leq G_{Q}(q_{1},q_{2})\big\}+O_{\varepsilon}(q_{1}^{1/2+\varepsilon})\bigg)$
$=\frac{1}{\zeta(2)}\int_{QI_{1}}\int_{QI_{2}}\frac{G_{Q}(u,v)-F_{Q}(u,v)}{u}\,dvdu+O_{\varepsilon}(Q^{3/2+\varepsilon})$
$=\frac{Q^{2}}{\zeta(2)}\iint_{\mathcal{R}}\frac{G_{Q}(Qx,Qy)-F_{Q}(Qx,Qy)}{Qx}\,dxdy+O_{\varepsilon}(Q^{3/2+\varepsilon})$
$=\frac{Q^{2}}{\zeta(2)}\iint_{\mathcal{R}}H_{2}(x,y)\,dxdy+O_{\varepsilon}(Q^{3/2+\varepsilon}),$
where
$H_{2}(x,y)=\min\bigg\{\frac{3x+2y-1}{x+y},1\bigg\}-\max\bigg\{\frac{3x-y-1}{x},\frac{x+(2-\lfloor x/y\rfloor)y-1}{y},0\bigg\},$ (13)
with $H_{2}(x,y)>0$ on $\mathring{\mathcal{V}}_{2}$.
Assume next that $\mathcal{R}\subseteq\mathring{\mathcal{V}}_{3}$, so $x_{0}+y_{0}<1$, $x_{0}<y_{0}$ and $3x_{0}+2y_{0}>1$. Here, the integer pairs $(q_{1},q_{2})\in\mathscr{D}_{Q}\cap Q\mathcal{R}$ must satisfy the inequalities $q_{1}+q_{2}<Q$, $q_{1}<q_{2}$ and $3q_{1}+2q_{2}>Q$. Taking (11) into account, this is equivalent with counting integer triples $(q_{1},q_{2},a)$ such that  [^3] Here, $a$ plays the role of $\bar{\bar{q}}_{1}$.
$\begin{split}&q_{1}\in QI_{1},\quad q_{2}\in QI_{2},\quad q_{1}<q_{2},\quad 0<a<q_{2},\quad q_{1}a\equiv 1\left(\mathrm{mod\;}q_{2}\right),\\ &\bigg(2+\bigg\lfloor\frac{q_{2}}{q_{1}}\bigg\rfloor\bigg)q_{1}-q_{2}+\frac{q_{1}a-1}{q_{2}}\leq Q,\quad(\Longleftrightarrow h(a_{1}/q_{1})\leq Q)\\ &q_{1}+q_{2}+a\leq Q,\quad(\Longleftrightarrow h(a_{2}/q_{2})\leq Q)\\ &2q_{1}+q_{2}+a+\frac{q_{1}a-1}{q_{2}}>Q.\quad(\Longleftrightarrow h(a_{*}/q_{*})>Q)\end{split}$ (14)
@@
Employing the same technique as in $\mathcal{R}\subseteq\mathring{\mathcal{V}}_{1}\cup\mathring{\mathcal{V}}_{2}$, the number of triples $(q_{1},q_{2},a)$ as in (14) is seen to be given by
$\sum\limits_{q_{2}\in QI_{2}}$ $\sum\limits_{\begin{subarray}{c}q_{1}\in QI_{1}\\ F_{Q}(q_{1},q_{2})\leq a\leq G_{Q}(q_{1},q_{2})\\ q_{1}a\equiv 1\left(\mathrm{mod\;}q_{2}\right)\end{subarray}}1$
$=\sum\limits_{q_{2}\in QI_{2}}\bigg(\frac{\varphi(q_{2})}{q_{2}^{2}}\operatorname{Area}\big\{(a,q_{1}):q_{1}\in QI_{1},F_{Q}(q_{1},q_{2})\leq a\leq G_{Q}(q_{1},q_{2})\big\}+O_{\varepsilon}(q_{2}^{1/2+\varepsilon})\bigg)$
$=\frac{1}{\zeta(2)}\int_{QI_{2}}\int_{QI_{1}}\frac{G_{Q}(u,v)-F_{Q}(u,v)}{v}\,dudv+O_{\varepsilon}(Q^{3/2+\varepsilon})$
$=\frac{Q^{2}}{\zeta(2)}\iint_{\mathcal{R}}\frac{G_{Q}(Qx,Qy)-F_{Q}(Qx,Qy)}{Qy}\,dxdy+O_{\varepsilon}(Q^{3/2+\varepsilon})$
$=\frac{Q^{2}}{\zeta(2)}\iint_{\mathcal{R}}H_{3}(x,y)\,dxdy+O_{\varepsilon}(Q^{3/2+\varepsilon}),$
@@
$H_{3}(x,y)=\min\bigg\{\frac{1+y-(2+\lfloor y/x\rfloor)x}{x},\frac{1-x-y}{y},1\bigg\}-\max\bigg\{\frac{1-2x-y}{x+y},0\bigg\},$ (15)
with $H_{3}(x,y)>0$ on $\mathring{\mathcal{V}}_{3}$. This completes the proof of Theorem 2.
@@ 4. A partition of the region $\mathcal{V}$
$\bigg(\frac{q_{1}}{Q},\frac{q_{2}}{Q}\bigg)\in\mathcal{W}_{1}:=\mathcal{V}_{1}.$
@@
$\bigg(\frac{q_{1}}{Q},\frac{q_{4}}{Q}\bigg)\in\mathcal{W}_{3,2}:=\bigg\{(x,y)\in[0,1]^{2}:\max\bigg\{1-2x,\frac{1-x}{3},2x-1\bigg\}\leq y\leq\frac{1-x}{2}\bigg\},$
with $\mathcal{W}_{3,2}\subseteq[1/3,3/5]\times[1/7,1/3]$.
@@
$\bigg(\frac{q_{1}}{Q},\frac{q_{4}}{Q}\bigg)\in\mathcal{W}_{3,3}:=\bigg\{(x,y)\in[0,1]^{2}:\max\bigg\{1-3x,\frac{1-x}{2}\bigg\}\leq y\leq 1-2x\bigg\},$
with $\mathcal{W}_{3,3}\subseteq[0,1/3]\times[1/3,1]$.
@@
$\bigg(\frac{q_{1}}{Q},\frac{q_{5}}{Q}\bigg)\in\mathcal{W}_{4,2}:=\bigg\{(x,y)\in[0,1]^{2}:\max\bigg\{1-2x,\frac{1-x}{4},2x-1\bigg\}\leq y\leq\frac{1-x}{3}\bigg\},$
with $\mathcal{W}_{4,2}\subseteq[2/5,4/7]\times[1/9,1/5]$.
@@
$\bigg(\frac{q_{1}}{Q},\frac{q_{5}}{Q}\bigg)\in\mathcal{W}_{4,3}:=\bigg\{(x,y)\in[0,1]^{2}:\max\bigg\{1-3x,\frac{1-x}{3}\bigg\}\leq y\leq\min\bigg\{\frac{1-x}{2},1-2x\bigg\}\bigg\},$
with $\mathcal{W}_{4,3}\subseteq[1/5,2/5]\times[1/5,2/5]$.
@@
$\bigg(\frac{q_{1}}{Q},\frac{q_{5}}{Q}\bigg)\in\mathcal{W}_{4,4}:=\bigg\{(x,y)\in[0,1]^{2}:\max\bigg\{1-4x,\frac{1-x}{2}\bigg\}\leq y\leq 1-3x\bigg\},$
with $\mathcal{W}_{4,4}\subseteq[0,1/5]\times[2/5,1]$.
@@
$\bigg(\frac{q_{1}}{Q},\frac{q_{6}}{Q}\bigg)\in\mathcal{W}_{5,2}:=\bigg\{(x,y)\in[0,1]^{2}:\max\bigg\{1-2x,\frac{1-x}{5},2x-1\bigg\}\leq y\leq\frac{1-x}{4}\bigg\},$
with $\mathcal{W}_{5,2}\subseteq[3/7,6/11]\times[1/11,1/7]$.
@@
$\bigg(\frac{q_{1}}{Q},\frac{q_{6}}{Q}\bigg)\in\mathcal{W}_{5,3}:=\bigg\{(x,y)\in[0,1]^{2}:\max\bigg\{1-3x,\frac{1-x}{4}\bigg\}\leq y\leq\min\bigg\{\frac{1-x}{3},1-2x\bigg\}\bigg\},$
with $\mathcal{W}_{5,3}\subseteq[1/4,3/7]\times[1/7,1/4]$.
@@
$\bigg(\frac{q_{1}}{Q},\frac{q_{6}}{Q}\bigg)\in\mathcal{W}_{5,4}:=\bigg\{(x,y)\in[0,1]^{2}:\max\bigg\{1-4x,\frac{1-x}{3}\bigg\}\leq y\leq\min\bigg\{\frac{1-x}{2},1-3x\bigg\}\bigg\},$
with $\mathcal{W}_{5,4}\subseteq[1/7,1/4]\times[1/4,3/7]$.
@@
$\bigg(\frac{q_{1}}{Q},\frac{q_{6}}{Q}\bigg)\in\mathcal{W}_{5,5}:=\bigg\{(x,y)\in[0,1]^{2}:\max\bigg\{1-5x,\frac{1-x}{2}\bigg\}\leq y\leq 1-4x\bigg\},$
with $\mathcal{W}_{5,5}\subseteq[0,1/7]\times[3/7,1]$.
Next, we show an example where the boundary line $3x+2y=1$ of $\mathcal{V}$ occurs as a side of a region $\mathcal{W}_{r,i}$. The tuple $(q_{1},\ldots,q_{10})=(27,95,68,41,96,55,69,83,97,14)$ represents consecutive elements in $\mathcal{F}_{100}$ with $q_{1}=27$ and $q_{10}=14$ consecutive in $\mathscr{S}_{100}$ and mediant $q_{1}+q_{10}=41=q_{4}$. The equalities $q_{2}=3q_{1}+q_{10}$, $q_{3}=2q_{1}+q_{10}$, $q_{4}=q_{1}+q_{10}$, $q_{5}=2q_{1}+3q_{10}$, $q_{6}=q_{1}+2q_{10}$, $q_{7}=q_{1}+3q_{10}$, $q_{8}=q_{1}+4q_{10}$, $q_{9}=q_{1}+5q_{10}$ show that the region $Q\mathcal{W}_{9,4}$ is defined by the inequalities $\max\{3q_{1}+q_{10},2q_{1}+3q_{10},q_{1}+5q_{10}\}\leq Q<\min\{q_{1}+q_{2}=4q_{1}+q_{10},q_{3}+q_{4}=3q_{1}+2q_{10},q_{6}+q_{7}=2q_{1}+5q_{10},q_{9}+q_{10}=q_{1}+6q_{10}\}$, and thus
@@
The region $\mathcal{W}_{9,4}$ is illustrated in the right-hand picture from Figure 2.